# Natural Language Processing for Dialects of a Language: A Survey

### v2

@@ 2.4. Recent work
One observes a renewed interest in using dialects to inform NLP tasks, as shown in Figure 1. The figure was generated using the ACL anthology (https://aclanthology.org/info/development/; Accessed on 9th January, 2024.). For “dialects”, we use ‘dialect’, ‘national variet’ (subword for inflections of ‘variety’), ‘national variation’, ‘creole’. For “socio-cultural”, we use the words “cultural”, and “socio-cultural”. We restrict to the year range 2000-2023.  (Hovy and Yang, 2021) shows that incorporating dialectic aspects is closely related to social factors of language. As a result, incorporating an understanding of dialects of a dataset is pivoting for fairer NLP tools.
Dialect awareness has been shown to improve the performance of NLP tasks such as machine translation (Sun et al., 2023), speech recognition (Plüss et al., 2023). Recent works have also focused on dialect-aware NLP tasks as in the case of machine translation of dialect to standard language translation as in the case of Chinese (Lu et al., 2022) (for Hokkein, a dialect of Chinese).
@@ 3. Scope & Trends
The focus of this survey is NLP approaches that are aware of dialects: either in the form of the choice of the dataset, incorporation in the model or evaluation along dimensions involving dialect. The survey provides a broad introduction to past NLP research on dialects spoken in different parts of the world. In the forthcoming subsections, we clarify the scope of this paper (Section 3.1) and highlight key trends (Section 3.2) that are described in detail in the following sections.
@@ 3.1. Scope
    Code-mixing: Code-mixing involves the use of words from two or more languages, often to reduce cognitive load. While code-mixing focuses on vocabulary, dialects are a combination of syntax and vocabulary.
@@
    Implicit selection biases: We also acknowledge that selection biases in datasets may introduce dialectic variations. For example, a dataset of tweets downloaded from a specific country is likely to have predominant dialects spoken in the country. However, we cannot locate these papers in particular, or, for social implications, claim that they are based on dialectic variations of a language without the authors mentioning so.
@@
    Accent variations: Finally, we focus on ‘text’-based research while acknowledging that the speech processing community has a rich history of using acoustic data centered around accent. This distinction between dialect and accent has been acknowledged in (Haugen, 1966). The focus on dialects allows us to restrict to the textual form of language which is the typical purview of NLP.
@@
    Systematic review: This survey is not a systematic review. We select key representative papers based on our interpretation of the innovation. We acknowledge that we may have missed out on important papers in the field. We will incorporate these papers as communicated by readers/reviewers. However, we cover a broad range of approaches in the survey.
@@ 3.2. Trends
    Tasks in focus: Older research dealt with dialectic datasets primarily for dialect classification. Past work shows performance degradation when the text contains dialects of a language as compared to the predominant (i.e., standard) form.
@@
    Languages in focus: The papers reporting work on dialects of Arabic are significantly higher than those for dialects of other languages. This has also been accelerated by research forums focussing on Arabic NLP. While the work in English is predominantly for African-American dialect of English, recent papers examine other dialects such as Indian English, Singaporean English and so on.
@@
    Mitigation is more than perturbation: Modifying a sentence or its representation to or from its dialectic variations has been achieved by perturbation techniques of varying complexity. However, recent papers show that dialect mitigation can be integrated into the model architecture itself using adversarial networks (Ball-Burack et al., 2021), hypernetworks (Xiao et al., 2023), etc.
@@
    Training resources: Labeled datasets, treebanks and lexicons in dialects of a language have been reported in the past. This includes datasets with dialect labels along with additional task-specific labels, where the task is an NLP research problem.
@@
    Evaluation datasets: NLP techniques evaluated on datasets in dialects have peculiar observations. Language identification classifiers produce lower performance when the text is in a dialect of a language. The performance of LLMs on dialectic datasets is positively correlated with socio-economic factors.
Figure 2 shows an overview of the approaches in terms of NLP for dialects. There have been different approaches to create labeled datasets, tree-banks and lexicons. In terms of models, past work varies in terms of NLP tasks and the way dialectic adaptation is handled: dialect transformation (where data is translated between dialects for the purpose of processing), dialect invariance (where models are made invariant to dialects) and dialect awareness (where models include dialect-specific components). Finally, we also describe dialectic datasets and resultant evaluations on downstream tasks including applications such as health monitoring. The next sections are centered around resources, NLU and NLG tasks keeping in view these aspects.
@@ 4. Resources
Being a data-driven field, NLP techniques rely on resources such as lexicons and textual datasets. In this section, we describe ways in which dialectic datasets have been created.
4.1. Dialectic Lexicons
Dialectic lexicons correspond to word lists or word mappings about a dialect. Past work in the creation of dialectic lexicons lies in three categories: the use of online dictionaries, and the use of textual corpora.
@@ 4.1.1. Online dictionaries
(Azouaou and Guellil, 2017) create a lexicon of words mapping French and its Algerian dialect. They use online dictionaries along with a combination of manual and automatic methods to enhance the lexicon. This includes many-to-one mapping of words in the two sets. Similarly,  (Boujelbane et al., 2013) build bilingual lexicons to create Tunisian dialectic corpora to adapt n-gram models for statistical machine translation.
@@ 4.1.2. Textual corpora
(Abdul-Mageed and Diab, 2014) present a lexicon of words in dialects of Arabic including Levantine and Egyptian dialects. They present two lexicons: adjectives from news articles and common words in online chat forums. The words are labeled with a combination of manual and automatic techniques, the latter based on statistical techniques such as pointwise mutual information. Similarly, (Burghardt et al., 2016) use the web as a corpus to create a lexicon for the Bavarian dialect of German. Starting with a corpus of Facebook comments, they provide a rule-based algorithm to create the lexicon. They first extract unique text forms, and then filter non-dialect words based on the Dortmund chat corpus.  (Younes et al., 2020; Harrat et al., 2018) discuss various existing resources for the Maghrebi Arabic dialects (MAD) including annotated corpora for language identification, and morpho-syntactic analysis. MAD include principally Algerian Arabic, Moroccan Arabic and Tunisian Arabic.
4.2. Dialectic Datasets
Figure 3. Overview of dataset creation approaches.
Figure 3 summarises dialectic datasets in terms of three dimensions: dialects, labeling techniques and data sources. Several data sources such as social media, conversation transcripts and so on have been used. Similarly, past work has also dealt with dialects of languages spoken around the world. In terms of procuring and labeling these datasets, the following methods have been used:
@@ 4.2.1. Recruit native speakers of specific language varieties
(Estival et al., 2014) create a dataset of audio-visual recordings of 1000 speakers of Australian English. The dataset is accompanied by a transcript which was manually created for 100 speakers.  (Bouamor et al., 2018b) present the MADAR: a manually curated parallel corpus of sentences in Arabic dialects along with English, French and Modern Standard Arabic. 131313 (Obeid et al., 2019) is a demonstration based on the dataset. Similarly, (Riley et al., 2023) create a parallel corpus of English sentences and two dialects each of Portuguese and Chinese with the help of native speakers of these dialects.  (Eisenstein et al., 2023) present a dataset of conversations between speakers playing the game of taboo. The dataset consists of speech recordings as well as text transcripts. Seddah et al. (2020) focus on treebank creation for Algerian supplemented with monolingual data obtained from CommonCrawl. They highlight the inherent difficulty of finding annotators and the cost of the same indicating the challenges for dialectic data generation. Riabi et al. (2023) further extend this with additonal layers of morpho-syntactic knowledge and correct errors in the same.
@@ 4.2.2. Perturbation
(Ziems et al., 2022) evaluate natural language understanding for African-American English. They design rules to perturb the dataset from Standard American English to African-American English. They then get them validated by native speakers. (Ziems et al., 2023) present Multi-VALUE, a suite of resources to evaluate fairness of LLMs by creating dialectic variations of a dataset. The suite provides mechanisms to generate 50 dialects of English by applying a set of perturbations.
@@ 4.2.3. Keywords
(Wang et al., 2017) create a dataset of Singaporean English sentences by searching for typical Singaporean English terms in online forums.  (Ramponi and Casula, 2023a) take a complementary approach to create a dataset of tweets in dialects of Italian along with other languages spoken in Italy (which are not necessarily derived from Italian). They use location-based search to obtain the set of tweets from different regions of interest from within Italy. Following this, they use out-of-vocabulary words to identify words that are indicative of geographical regions and, as a result, dialects. A related dataset is GeoLingIt (Ramponi and Casula, 2023b). In the case of social media, hashtags can be used to obtain datasets in certain dialects.  (Scherrer et al., 2023) take advantage of dialect awareness week in Finland. They use a hashtag indicating usage of dialects in order to collect tweets in different dialects of Finnish.
@@ 4.2.4. Location
: Data from particular geographics can be extracted using filters (where the location is known) or inference (where it is not known).  (Jurgens et al., 2017) use location-based filters available on Twitter at the time. They use language identification classifiers to predict the language and identify dialectic users. (Husain et al., 2022) obtain tweets from Kuwait to create a dataset of tweets in the Kuwaiti dialect of Arabic.  (Coats, 2022) create an unlabeled dataset of Youtube comments. They start with a list of councils in Australia, extract official Youtube channels and retrieve comments. They manually validate the correctness of the channels. When working with geographically dispersed dialects, sampling may also be used.  (Dunn and Adams, 2020) create a Web-based corpus in different dialects of the corpus by sampling sentences from different countries. The goal is to build a Web-based corpora where the number of instances are reflective of the population of speakers in a country. The paper states that such a geography-aware corpus can lead to geography-aware representations when language models are trained on them. Some researchers perform location inference using computational techniques.  (Blodgett et al., 2018) use a demographically-aligned topic model that predicts a demographic for a tweet. They use it as a label followed by manual validation.
@@ 4.2.5. Dialect-aware annotation
: One such example is by (Sap et al., 2019). They examine racial bias towards African-American English in the case of hate speech detection. They propose race and dialect priming in order to improve the quality of annotation. In order to prime the annotators, they propose to ask two questions: (a) is the tweet offensive to them, and (b) is the tweet offensive to anyone. The dialect and race of the speaker are shown to the annotators.
Several datasets exist for varieties of Arabic, including Palestinian Arabic (Jarrar et al., 2017; Dibas et al., 2022), Gulf Arabic (Khalifa et al., 2016), Egyptian Arabic (Maamouri et al., 2014), and Bahraini Arabic (Abdulrahim et al., 2022). This is in stark contrast with the lack of availability of datasets for dialects of English or several other languages of the world.
Paper Source languages
(Farha and Magdy, 2022) Dialects of Arabic
(Lui and Cook, 2013) Dialects of English
(Simaki et al., 2017) Dialects of English
(Hanani and Naser, 2020) Dialect of Arabic
(Goswami et al., 2020) Dialect of German
(Hou and Huang, 2020) Dialects of Chinese
(Abdul-Mageed et al., 2018) Dialects of Arabic
(Blodgett et al., 2016) Dialect of English
(Barnes et al., 2021) Dialects of Norwegian
(Aji et al., 2022) Dialects of Indonesian
(Ramponi and Casula, 2023a) Dialects of Italian
Table 3. Language distribution for dialect detection.
@@ 5. Natural Language Understanding (NLU)
This section covers NLU approaches centered around dialects. This includes approaches for NLP tasks such as dialect identification, sentiment analysis, morphosyntactic analysis and parsing. We also describe approaches reported on NLU benchmarks which cover multiple tasks.
@@ 5.1. Dialect Identification
The most commonly researched task in the scope of this paper is dialect identification. Dialect identification deals with the prediction of the dialect of an input text. Early approaches to dialect identification employed distance-based metrics, namely, Levenshtein, Manhattan, and Euclidean distance with clustering and training a Kohonen map  [^14] A Kohonen map is a neural network in which concepts are plotted during a training procedure. Similar concepts are plotted near one another (Nerbonne and Heeringa, 1997, 2002). They indicate that feature representations are more sensitive, and that, Manhattan distance and Euclidean distance are good measures of phonetic overlap. (Elnagar et al., 2021) is a systematic review of identification of dialects of Arabic. For dialects of Arabic, lexical resources such as lexicons and tree banks, and models using SVM or sequential neural layers like BiLSTM have been reported. (Jauhiainen et al., 2019) is a survey of automatic language identification. They mention that dialect detection may be more difficult than language detection since dialects may have lexical or syntactic overlap. In doing so, the survey does not make a distinction between languages and dialects - and treats different dialects as different class labels, while still maintaining a classification approach. However, one sees challenges in this regard. (Blodgett et al., 2016) use an extension of an LDA-based model for language identification. The language identification classifier is expected to predict dialects of English as English. However, the results show that the classifier trained on ‘standard English’ data may not work as well for dialectic variations: as seen in the case of African-American English. (Boujou et al., 2021) benchmark a novel dataset of  50,000 tweets for five dialects of Arabic- Algerian, Lebanon, Morocco, Tunisian, and Egyptian, and also a present baseline benchmark which utilizes classical machine learning. While the majority of past work defines dialect identification as a Boolean/multi-class classification,  (Baimukan et al., 2022) use a hierarchy of dialect labels based on geographical and linguistic proximity.
We now describe details of past work in dialect identification in terms of shared tasks, datasets, and pre-deep learning and deep learning-based approaches.

### v4

@@ 2.4. Recent work
One observes a renewed interest in using dialects to inform NLP tasks, as shown in Figure 2. The figure was generated using the ACL anthology [^2] https://aclanthology.org/info/development/; Accessed on 9th January, 2024.. For “dialects”, we use ‘dialect’, ‘national variety’ (subword for inflections of ‘variety’), ‘national variation’, and ‘Creole’. For “socio-cultural”, we use the words “cultural”, and “socio-cultural”. We restrict to the year range 2000-2023.
Dialect awareness has been shown to improve the performance of NLP tasks such as machine translation (Sun et al., 2023), speech recognition (Plüss et al., 2023). Recent works have also focused on dialect-aware NLP tasks as in the case of machine translation of dialect to standard language translation as in the case of Chinese (Lu et al., 2022) (for Hokkien, a dialect of Chinese).
@@ 3. Scope & Trends
The focus of this survey is on NLP approaches that are aware of dialects: either in the form of the choice of the dataset, incorporation in the model or evaluation along dimensions involving dialect. The survey provides a broad introduction to past NLP research on dialects spoken in different parts of the world. In the forthcoming subsections, we clarify the scope of this paper (Section 3.1) and highlight key trends (Section 3.2) that are described in detail in the following sections.
@@ 3.1. Scope
    Code-mixing: Code-mixing involves the use of words from two or more languages, often to reduce cognitive load. This survey does not focus on code-mixing.
@@
    Implicit selection biases: We also acknowledge that selection biases in datasets may introduce dialectal variations. For example, a dataset of tweets downloaded from a specific country is likely to have predominant dialects spoken in the country. However, we cannot locate these papers in particular, or, for social implications, claim that they are based on dialectal variations of a language without the authors mentioning so.
@@
    Accent variations: Finally, we focus on ‘text’-based research while acknowledging that the speech processing community has a rich history of using acoustic data centered around accent. To this end, we briefly touch upon speech since dialects and speech are intertwined to a certain degree. However, our primary focus is on text since the text aspect has received a lot more attention than the speech aspect. This also sets up a situation where a future survey can expand on the speech aspect of dialectal processing. The focus on the textual form is the typical purview of NLP.
@@
    Systematic review: This survey is not a systematic review in the sense that we do not exhaustively cover all works on dialects due to limited time and paper space. Instead, we select key representative papers based on our interpretation of the innovation, which, according to us, cover key progress and innovation in NLP for dialects. We acknowledge that we may have missed out on some important papers in the field. We will incorporate these papers as communicated by readers/reviewers. However, we cover a broad range of approaches in the survey.
@@ 3.2. Trends
    Tasks in focus: Older research dealt with dialectal datasets primarily for dialect classification. Past work shows performance degradation when the text contains dialects of a language as compared to the predominant (i.e., standard) form.
@@
    Languages in focus: The papers reporting work on dialects of Arabic are significantly more than those for dialects of other languages. This has also been accelerated by research forums focusing on Arabic NLP. While the work in English is predominantly for the African-American dialect of English, recent papers examine other dialects such as Indian English, Singaporean English and so on.
@@
    Mitigation is more than perturbation: Modifying a sentence or its representation to or from its dialectal variations has been achieved by perturbation techniques of varying complexity. However, recent papers show that dialect mitigation can be integrated into the model architecture itself using adversarial networks (Ball-Burack et al., 2021), hypernetworks (Xiao et al., 2023), etc.
@@
    Training resources: Labeled datasets (including treebanks) and lexicons in dialects of a language have been reported in the past. This includes datasets with dialect labels along with additional task-specific labels, where the task is an NLP research problem.
@@
    Evaluation datasets: NLP techniques evaluated on datasets in dialects have peculiar observations. Language identification classifiers produce lower performance when the text is in a dialect of a language. The performance of LLMs on dialectal datasets is positively correlated with socio-economic factors.
Figure 3 shows an overview of the approaches in terms of NLP for dialects. There have been different approaches to create labeled datasets, tree-banks and lexicons. In terms of models, past work varies in terms of NLP tasks and the way dialectal adaptation is handled: dialect transformation (where data is translated between dialects for the purpose of processing), dialect invariance (where models are made invariant to dialects) and dialect awareness (where models include dialect-specific components). Finally, we also describe dialectal datasets and resultant evaluations on downstream tasks including applications such as health monitoring.
@@ 4. Resources
Being a data-driven field, NLP techniques rely on resources such as lexicons and textual datasets. In this section, we describe ways in which dialectal datasets have been created.
4.1. Dialectal Lexicons
Dialectal lexicons correspond to word lists or word mappings about a dialect. Although lexicons were popular in early approaches of NLP, a recent paper by  Artemova and Plank (2023) highlights the potential of dialectal lexicons and describes an approach to create such lexicons using large language models. Prior to this, research in the creation of dialectal lexicons lies in three categories: the use of online dictionaries, and the use of textual corpora.
@@ 4.1.1. Online dictionaries
Azouaou and Guellil (2017) create a lexicon of words mapping French and its Algerian dialect. They use online dictionaries along with a combination of manual and automatic methods to enhance the lexicon. This includes many-to-one mapping of words in the two sets. Similarly,  Boujelbane et al. (2013) build bilingual lexicons to create Tunisian dialectal corpora to adapt n-gram models for statistical machine translation.
@@ 4.1.2. Textual corpora
Abdul-Mageed and Diab (2014) present a lexicon of words in dialects of Arabic including Levantine and Egyptian dialects. They present two lexicons: adjectives in news articles and common words in online chat forums. The words are labeled with a combination of manual and automatic techniques, the latter based on statistical techniques such as pointwise mutual information. Similarly, Burghardt et al. (2016) use the web as a corpus to create a lexicon for the Bavarian dialect of German. Starting with a corpus of Facebook comments, they provide a rule-based algorithm to create the lexicon. They first extract unique text forms, and then filter non-dialect words based on the Dortmund chat corpus.  Younes et al. (2020); Harrat et al. (2018) discuss various existing resources for the Maghrebi Arabic dialects (MAD) including annotated corpora for language identification, and morpho-syntactic analysis. MAD include principally Algerian Arabic, Moroccan Arabic and Tunisian Arabic.
4.1.3. Lexicon Induction Using LLMs
A recent approach by  Artemova and Plank (2023) performs German dialect lexicon induction using LLMs.
4.2. Dialectal Datasets
Datasets based on different data sources (such as social media, and conversation transcripts) and dialects have been reported. In terms of procuring and labeling these datasets, the following methods have been used:
@@ 4.2.1. Recruit native speakers of specific language varieties
Estival et al. (2014) create a dataset of audio-visual recordings of 1000 speakers of Australian English. The dataset is accompanied by a transcript, which was manually created for 100 speakers.  Bouamor et al. (2018) present MADAR: a manually curated parallel corpus of sentences in Arabic dialects along with English, French and Modern Standard Arabic. 444 Obeid et al. (2019) is a demonstration based on the dataset. Similarly, Riley et al. (2023) create a parallel corpus of English sentences and two dialects each of Portuguese and Chinese with the help of native speakers of these dialects.  Eisenstein et al. (2023) introduce MD-3 a dataset of conversations between speakers playing the game of taboo. The dataset consists of speech recordings as well as text transcripts. Seddah et al. (2020) focus on treebank creation for Algerian supplemented with monolingual data obtained from CommonCrawl. They highlight the inherent difficulty of finding annotators and the cost of the same indicating the challenges for dialectal data generation. Riabi et al. (2023) further extend this with additonal layers of morpho-syntactic knowledge and correct errors in the same.
@@ 4.2.2. Perturbation
Ziems et al. (2022) evaluate natural language understanding for African-American English. They design rules to perturb the dataset from Standard American English to African-American English. They then get them validated by native speakers. Ziems et al. (2023) present Multi-VALUE, a suite of resources to evaluate fairness of LLMs by creating dialectal variations of a dataset. The suite provides mechanisms to generate 50 dialects of English by applying a set of perturbations. Messner and Lippincott (2024) present a dataset of 19th century American literary orthovariant tokens with a novel layer of human-annotated dialect group tags, to examine language modelling assumptions, and find evidence that choice of tokenization scheme meaningfully impact the type of orthographic information in a language model.
@@ 4.2.3. Keywords
Wang et al. (2017) create a dataset of Singaporean English sentences by searching for typical Singaporean English terms in online forums.  Ramponi and Casula (2023a) take a complementary approach to create a dataset of tweets in dialects of Italian along with other languages spoken in Italy (which are not necessarily derived from Italian). They use location-based search to obtain the set of tweets from different regions of interest from within Italy. Following this, they use out-of-vocabulary words to identify words that are indicative of geographical regions and, as a result, dialects. A related dataset is GeoLingIt (Ramponi and Casula, 2023b). In the case of social media, hashtags can be used to obtain datasets in certain dialects.  Kuparinen (2023) take advantage of dialect awareness week in Finland. They use a hashtag indicating usage of dialects in order to collect tweets in different dialects of Finnish. In the context of Arabic dialect tweets,  Boujou et al. (2021) benchmark is a novel dataset of  50,000 tweets for five dialects of Arabic-Algerian, Lebanon, Morocco, Tunisian, and Egyptian.
@@ 4.2.4. Location
Data from particular geographics can be extracted using filters (where the location is known) or inference (where it is not known).  Jurgens et al. (2017) use location-based filters available on Twitter at the time. They use language identification classifiers to predict the language and identify dialectal users. Husain et al. (2022) obtain tweets from Kuwait to create a dataset of tweets in the Kuwaiti dialect of Arabic.  Coats (2022) create an unlabeled dataset of Youtube comments. They start with a list of councils in Australia, extract official Youtube channels and retrieve comments. They manually validate the correctness of the channels. When working with geographically dispersed dialects, sampling may also be used. Hovy and Purschke (2018) use Doc2Vec on a large corpus of anonymous online posts to learn document representation of cities, and recover dialect areas using geographic information via retrofitting and agglomerative clustering. Dunn and Adams (2020) create a Web-based corpus in different dialects by sampling sentences from different countries. The goal is to build a Web-based corpus where the number of instances is reflective of the population of speakers in a country. The paper states that such a geography-aware corpus can lead to geography-aware representations when language models are trained on them. A criticism to the location-based filtering mechanism is by  Goutte et al. (2016).
@@ 4.2.5. Dialect-aware annotation
One such example is by Sap et al. (2019). They examine racial bias towards African-American English in the case of hate speech detection. They propose race and dialect priming in order to improve the quality of annotation. In order to prime the annotators, they propose to ask two questions: (a) is the tweet offensive to them?, and (b) is the tweet offensive to anyone? The dialect and race of the speaker are shown to the annotators.
Several datasets exist for varieties of Arabic (Diab et al., 2010), including Palestinian Arabic (Jarrar et al., 2017; Dibas et al., 2022), Gulf Arabic (Khalifa et al., 2016), Egyptian Arabic (Maamouri et al., 2014), and Bahraini Arabic (Abdulrahim et al., 2022). This is in stark contrast with the lack of availability of datasets for dialects of English or several other languages of the world.
@@ 5. Natural Language Understanding (NLU)
This section covers NLU approaches centered around dialects. This includes approaches for NLP tasks such as dialect identification, sentiment analysis, morphosyntactic analysis and parsing. We also describe approaches reported on NLU benchmarks, which cover multiple tasks.
@@ 5.1. Dialect Identification
The most commonly researched task in the scope of this paper is dialect identification. Dialect identification deals with the prediction of the dialect of an input text. Early approaches to dialect identification employed distance-based metrics, namely, Levenshtein, Manhattan, and Euclidean distance with different clustering techniques (Nerbonne and Heeringa, 1997, 2002). They indicate that feature representations are more sensitive, and that Manhattan distance and Euclidean distance are good measures of phonetic overlap. Elnagar et al. (2021) is a systematic review of identification of dialects of Arabic. For dialects of Arabic, lexical resources such as lexicons and treebanks, and models using SVM or sequential neural layers like BiLSTM have been reported. Jauhiainen et al. (2019) is a survey of automatic language identification. They describe that dialect detection may be more difficult than language detection since dialects may have lexical or syntactic overlap. In doing so, the survey does not make a distinction between languages and dialects - and treats different dialects as different class labels, while still maintaining a classification approach. However, one sees challenges in this regard.  Boujou et al. (2021) present a baseline approach, which utilises classical machine learning. While the majority of past work defines dialect identification as a Boolean/multi-class classification,  Baimukan et al. (2022) use a hierarchy of dialect labels based on geographical and linguistic proximity. We now describe details of past work in dialect identification in terms of shared tasks, datasets, and pre-deep learning and deep learning-based approaches.
Shared Task Dialects/Languages
(Zampieri et al., 2014) Brazilian Portuguese and European Portuguese; American and British English; and Argentinian Spanish and Castilian Spanish
(Zampieri et al., 2015) American and British English; and Argentinian Spanish and Castilian Spanish
(Malmasi et al., 2016) Dialects of English, Spanish, French and Arabic
(Zampieri et al., 2019) Dialects of German, Chinese, Romanian
(Gaman et al., 2020) Dialects of Romanian, Geolocation-based Varieties
(Aepli et al., 2022) Dialects of French and Italian
(Aepli et al., 2023) Dialects of Indo-European and Ural languages (and other tasks)
(Abdul-Mageed et al., 2023) Dialects of Arabic
Table 3. Shared tasks related to dialect identification.